(6) For n=3, Verify the following summation formula: Σ (1) 2 3.
Answer #1

Wennow that (B) n! Y! (N-1)! 123 3 i 2 LHS (3) izo = ( )2+()+(3):2+ (3) 23 1. + 3.2 + 3.97 1.8 = 1+ 6 + 12 + 8 = = 27 R.HS 3If you don't understand please comment